please answer the following questions:

1) An island has 800 young dependents, 400 elderly dependents, and 200 people of working age, please calculate:
-Child dependency ratio:____(per 100 persons)
-Total dependency ratio:____(per 100 persons)

2)_____ is the proportion of a population that is elderly.
Answer #1

total population is 800+400+200=1400

1. Child Dependency ratio is (800/1400)*100=57.14

Total Dependency ratio is (1200/1400)*100= 85.71

2. (400/1400)*100=28.57